OpenStudy (anonymous):

An inequality is comparing two positive numbers. Explain why multiplying by a negative number on both sides of the inequality reverses the sign...

OpenStudy (anonymous):

suppose you have two positive numbers like \(3\) and \(5\) where clearly \(3<5\)

OpenStudy (anonymous):

if you multiply by a negative number, say \(-1\) then you have \(-3\) and \(-5\) so the inequality has switched \[3<5\\ -3>-5\]

OpenStudy (anonymous):

for a more mathematical explanation \[a<b\] means there is a positive number \(p\) where \(a+p=b\) if you multiply by a negative number, say \(-2\) you will have \[-2a-2p=-b\] making \[-2a=-b+2p\] and so \[-2a>-2b\]

OpenStudy (anonymous):

Thank you.